OpenStudy (anonymous):

Solve for x in the proportion 9/27x=6/x-1 . x = 2 and x = 0 x ≈ −2.48 and x ≈ 2.48 x ≈ −0.85 and x ≈ 1.18 x ≈ −6 and x ≈ 6.16

jimthompson5910 (jim_thompson5910):

The equation is \[\frac{9}{27x} = \frac{6}{x-1}\] right?

OpenStudy (anonymous):

Yes

OpenStudy (goldphenoix):

It should be. To solve the equation, cross multiply.

jimthompson5910 (jim_thompson5910):

\[\large \frac{9}{27x} = \frac{6}{x-1}\] turns into \[\large 9(x-1) = 6*27x\] after you cross multiply what's next?

OpenStudy (goldphenoix):

Oh. How do you make the font bigger? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

by sticking a \large at the front of the code

OpenStudy (anonymous):

@jim_thompson5910 Distibutive property?

OpenStudy (goldphenoix):

Oh thanks.

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (goldphenoix):

Yes, Caitlynn. You use the distributive property for the left side.

OpenStudy (anonymous):

So itll come out to 9x-9=6*27x? What do i do next?

OpenStudy (goldphenoix):

Simplify the right side.

OpenStudy (anonymous):

9x-9=162x?

OpenStudy (goldphenoix):

Correct.

OpenStudy (anonymous):

What would i do next?

OpenStudy (goldphenoix):

You can guess and check. :D

OpenStudy (anonymous):

Do i divide the 162x by 9x which would equal 18x So id have -9=18x. take the 18 divide it by -9 x=-2?

OpenStudy (goldphenoix):

Yes you can do that as well.

OpenStudy (anonymous):

wait thats not an answer... I'm confused now.

OpenStudy (goldphenoix):

Ask @jim_thompson5910 He's the expert.

OpenStudy (anonymous):

Okay.

jimthompson5910 (jim_thompson5910):

there has to be a typo in the original problem this is a linear equation with one solution (not 2)

jimthompson5910 (jim_thompson5910):

but each answer choice lists two answers

OpenStudy (goldphenoix):

I thought so too.

OpenStudy (anonymous):

Solve for x in the proportion . x = 2 and x = 0 x ≈ −2.48 and x ≈ 2.48 x ≈ −0.85 and x ≈ 1.18 x ≈ −6 and x ≈ 6.16

OpenStudy (goldphenoix):

That's the original problem?

OpenStudy (anonymous):

Yes.

OpenStudy (goldphenoix):

No wonder. You typed it wrong.

OpenStudy (goldphenoix):

\[\large\large\frac{ 9 }{ 27x } = \frac{ x }{ x+3 }\]

OpenStudy (goldphenoix):

So review what Jim and I taught you.

OpenStudy (goldphenoix):

I shall rewind this. You can cross multiply. After cross multiplying, it should end up: 9(x+3) = 27x(x) Use the distributive property on the left to get: 9x+27 Simplify the one in the right: \[\large27x^2\] Now your equation should look like: 9x+27 = 27x^2.

OpenStudy (anonymous):

okay i got that. so divide 27x^2 by 9x?

OpenStudy (anonymous):

so would that be 3x?

OpenStudy (goldphenoix):

Yes.

OpenStudy (anonymous):

after that i get 9. But thats not an answer so im stuck now

OpenStudy (goldphenoix):

After that, it should give you 27 = 3x.

OpenStudy (goldphenoix):

Hmm... Let me check my work.

OpenStudy (anonymous):

Okay.

OpenStudy (goldphenoix):

@jim_thompson5910 Check my work please. I think I did something wrong. T~T

OpenStudy (goldphenoix):

I used a calc. It's suppose to be C, but I don't know what I did wrong. =/

OpenStudy (anonymous):

oh. okay. its fine though.